OpenStudy (anonymous):

e[3.05] Given the function g(x) = −4x + 5, find g(−4). −11 21 2.25 11

OpenStudy (anonymous):

@Vackious

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

Substitute x = -4 into the equation

OpenStudy (akashdeepdeb):

Substitute the value in the EQUATION! g(a) = -4(a) + 5 Similarly get the ANSWER!! :)

OpenStudy (anonymous):

i dont get it @AkashdeepDeb

OpenStudy (akashdeepdeb):

Put -4 in g(x) -4(-4) + 5 SI THE ANSWER!! :) Get it now! :D

OpenStudy (anonymous):

how to put -4 in g(x)

OpenStudy (akashdeepdeb):

f(x) or g(x) or anything like that is a function! That means they will have one value for one x value ok? So it is given that g(x) = -4.(x) + 5 And we have to find g(-4) So just put g(-4) = -4.-4 + 5 [REPLACE X WITH -4] --> As simple as that!! :D And then calculate the final answer!! :)

OpenStudy (anonymous):

so 21?

OpenStudy (akashdeepdeb):

YES!! :)